# Bridging Privacy and Robustness for Trustworthy Machine Learning

### v4

@@ Abstract
The advent of machine learning has led to transformative changes across various domains, but the sensitive nature of data raises concerns about privacy and security. While Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary, such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP). This paper explores the intricate relationships between LDP, ABP, and MBP, and their implications for algorithmic robustness. We establish theoretical connections between these privacy notions, proving that LDP implies MBP and vice versa under certain conditions, and deriving bounds connecting MBP and ABP. We also investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms and construct privacy-preserving algorithms from PAC robust ones. Our findings provide valuable insights for constructing privacy-preserving and robust machine learning algorithms.
Index Terms:
Privacy-Preserving Algorithms; PAC Robust Learning; Trustworthy Machine Learning; Theoretical Connections
I Introduction
The advent of machine learning has heralded transformative changes across various domains, from healthcare to finance, by extracting meaningful insights from vast swathes of data. However, this data often contains sensitive information, raising legitimate concerns regarding privacy and security. The traditional paradigm of Local Differential Privacy (LDP) has been a cornerstone in addressing these concerns, aiming to ensure that the influence of any individual data point on the output of a computation is bounded, thereby obfuscating its presence or absence (Dwork, 2008). While LDP provides a strong foundation for privacy protection, it is essential to consider the adversary’s methods for potentially compromising privacy.
Recent research has proposed privacy concepts aligned with the Bayesian inference perspective of an adversary (He et al., 2019; Zhu et al., 2019), such as Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP). These concepts have emerged from an understanding that adversaries may possess prior knowledge and use their Bayesian reasoning capabilities to infer sensitive information from the output of machine learning algorithms. By incorporating the adversary’s perspective, these new paradigms offer a more comprehensive approach to privacy protection, taking into account the potential strategies an adversary might employ to infer sensitive information.
As these Bayesian privacy concepts gain traction, it is crucial to thoroughly investigate their relationships with traditional privacy measures like LDP. Understanding the connections and differences between these approaches will enable researchers and practitioners to develop more robust and effective privacy-preserving mechanisms. By leveraging the strengths of both LDP and Bayesian privacy concepts, we can work towards building machine learning systems that safeguard sensitive information while still extracting valuable insights from data.
In this paper, we aim to explore the intricate relationships between LDP, ABP, and MBP, providing a deeper understanding of how these privacy notions interact and complement each other. Through rigorous mathematical analysis and empirical studies, we seek to uncover the key factors that influence the effectiveness of these privacy protection mechanisms and provide guidance on how to design and implement them in real-world machine learning applications. By bridging the gap between traditional and Bayesian privacy concepts, we hope to contribute to the development of more secure and trustworthy machine learning systems that respect individual privacy while unlocking the full potential of data-driven insights. Through meticulous definitions and theorems, we elucidate how these privacy concepts are interconnected. A key result (Theorem IV.4) establishes that, under certain conditions, LDP implies MBP and vice versa. We also derive bounds connecting MBP and ABP (Theorem IV.6), providing valuable insights for constructing privacy-preserving algorithms.
The intersection of privacy and learning theory, particularly Probably Approximately Correct (PAC) learning, has attracted considerable scholarly inquiry. Groundbreaking research by Kasiviswanathan et al. (2011) revealed that a vast array of classes learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size. Building upon these foundational intersections, our work extends the discourse by delving into the robustness implications of privacy guarantees. Robustness in machine learning, the ability of algorithms to maintain performance despite inaccuracies or noise in the input data, has become increasingly important as models are deployed in unpredictable real-world environments. The concept of robustness has been traditionally studied within the PAC framework (Cullina et al., 2018), which provides a theoretical foundation for the performance of learning algorithms under perturbations. However, while previous studies have predominantly focused on sample complexity and learnability, the intersection of robustness with privacy has not been fully explored, particularly how robust learning can contribute to, or benefit from, privacy-preserving mechanisms. In this paper, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and conversely, how robust learning algorithms can inform the design of privacy-preserving mechanisms. We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from algorithms satisfying specific privacy standards (Theorem IV.7). This exploration aims to bridge the gap between robustness and privacy, providing valuable insights for developing machine learning algorithms that are both resilient to perturbations and secure against privacy breaches.
The main contributions of this paper are as follows:
  • We establish theoretical connections between different privacy notions, proving that LDP implies MBP and vice versa under certain conditions (Theorem IV.4) and deriving relationship between MBP and ABP (Theorem IV.6).
  • We investigate the relationship between PAC robust learning and privacy preservation, demonstrating how to derive PAC robustness from privacy-preserving algorithms (Theorem IV.7) and construct privacy-preserving algorithms from PAC robust ones.
@@ I Introduction
The remainder of this paper is organized as follows. Section II-A reviews related work on privacy measurements, Section II-D discusses robustness measurements, and Section II-E explores the relationship between privacy and robustness in the literature. In Section III, we introduce our privacy attack and defense framework. In Section IV, we define metrics for privacy and robustness, and establish theoretical connections between privacy notions and PAC robustness. We also introduce practical estimation approaches and analyses for MBP and ABP in F and G. Finally, we conclude the paper and discuss future research directions.
@@
TABLE I: Table of Notation
@@
II Related Work
II-A Privacy Measurements
Differential privacy (DP), a rigorously formalized privacy framework pioneered by Dwork et al. (2006), has been widely adopted to protect individual privacy in database systems. Abadi et al. (2016) proposed a method for training neural networks with a limited privacy budget, combining advanced machine learning methods and privacy-preserving mechanisms. This method is applicable to various optimization methods, including Momentum Rumelhart et al. (1986) and AdaGrad Duchi et al. (2011). Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), an adaptation of DP that incorporates data distribution considerations, along with a theoretical analysis and privacy computing framework, demonstrating its advantages through experimental results.
Rassouli and Gündüz (2019) employed average total variation distance (TV distance) for privacy measurement and minimum mean-square error (MMSE), mutual information, and error probability to assess utility, aiming to find the best balance between utility and privacy in data analysis tasks. Blauw and von Solms (2017) proposed two main privacy metrics based on an information-theoretic approach: information entropy and mutual information, to quantify and define online user privacy.
Achieving pure DP can be challenging in practical learning scenarios. Eilat et al. (2020) quantified the loss of privacy using relative entropy (Kullback-Leibler divergence) and defined the concept of privacy loss, referring to the information leakage caused by the posterior update of the agent’s prior beliefs. Foulds et al. (2016) discussed the constraints of obtaining privacy via posterior sampling and highlighted the benefits of the Laplace mechanism in Bayesian inference from both theoretical and empirical perspectives.
Other works have considered privacy measures akin to our approach. In social network settings, Alemany et al. (2019) proposed two metrics (Audience and Reachability) for the privacy risk of sharing information. du Pin Calmon and Fawaz (2012) discussed the minimization of maximum information leakage in privacy-preserving mapping, proposing a mathematical model and a formula that minimizes the maximum information leakage, intended to balance privacy against utility constraints, with privacy leakage assessed using KL divergence. In contrast, our work adopts Jensen-Shannon (JS) divergence for measuring privacy leakage, in line with Zhang et al. (2022, 2023d), due to its symmetric nature and the fact that its square root meets the triangle inequality Nielsen (2019); Endres and Schindelin (2003), aiding in quantifying privacy-utility trade-offs more effectively.
II-B Relation between Distinct Privacy Measurements
The advent of differential privacy has inspired numerous scholars to propose new privacy metrics and explore their relationships. Table III summarizes key relations between various privacy metrics, including our proposed Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP).
TABLE II: Relations between various privacy metrics.
@@ II-B Relation between Distinct Privacy Measurements
Dwork and Rothblum (2016) Concentrated DP DP $\Rightarrow$ Concentrated DP
Mironov (2017) Rényi-DP Rényi-DP $\Rightarrow$ $(\epsilon,\delta)$-DP
Mironov et al. (2009)2 IND-CDP, SIM-CDP SIM-CDP $\Rightarrow$ SIM∀∃-CDP3 $\Rightarrow$ IND-CDP
Dong et al. (2019) $f$-DP Gaussian DP4 $\Rightarrow$ Rényi-DP
Blum et al. (2008) Distributional Privacy Distributional Privacy $\Rightarrow$ DP
Triastcyn and Faltings (2020) Bayesian-DP Bayesian DP $\Rightarrow$ DP
Cuff and Yu (2016) Mutual-Information DP MI-DP $\Leftrightarrow$ DP (asymptotically)
Our Work $\xi$-MBP, $\epsilon$-ABP $\xi$-LDP $\Rightarrow$ $\xi$-MBP $\Rightarrow$ $2\xi$-LDP, MBP $\Rightarrow$ ABP
@@
    All "DP"s are shorthand for differential privacy.
@@
    This article proposes Computational Differential Privacy (CDP). SIM-CDP and IND-CDP are two concrete kinds of CDP.
@@
    Gaussian DP is a specific family of $f$-DP.
The introduction of differential privacy has catalyzed the development of numerous privacy metrics. Dwork and Rothblum (2016) proposed Concentrated Differential Privacy (Concentrated-DP), which relaxes differential privacy based on the mean privacy leakage of a randomized mechanism. Mironov (2017) introduced Rényi Differential Privacy (RDP), a variant based on Rényi divergence that unifies concentrated-DP, $\epsilon$-DP, and $(\epsilon,\delta)$-DP. Mironov et al. (2009) developed Computational Differential Privacy (CDP), which measures privacy against computationally-bounded adversaries. They introduced IND-CDP and SIM-CDP, establishing that SIM-CDP implies SIM∀∃-CDP, which in turn implies IND-CDP. Dong et al. (2019) presented f-Differential Privacy (f-DP) and Gaussian Differential Privacy as relaxations of traditional DP. Blum et al. (2008) proposed Distributional Privacy, offering stronger protection than traditional differential privacy. Triastcyn and Faltings (2020) introduced Bayesian Differential Privacy (BDP), which considers the dataset’s distribution and implies traditional DP.
Recent research has focused on creating taxonomies and comparing various privacy definitions. Cuff and Yu (2016) provided a comprehensive comparison of privacy notions, including differential privacy and its variants, introducing Mutual-Information Differential Privacy (MI-DP) and demonstrating its asymptotic equivalence to traditional differential privacy. Meiser (2018) examined the nuances of differential privacy, emphasizing the importance of distinguishing between approximate and probabilistic differential privacy. Sommer et al. (2019) proposed a unified approach to measuring privacy loss in sensitive data through privacy loss distribution (PLD), revealing that for non-adaptive mechanisms, privacy loss tends toward a Gaussian distribution, enabling tighter privacy bounds and more accurate calculations for mechanisms such as the Gaussian mechanism under ADP.
Desfontaines and Pejó (2019) classified privacy definitions into seven distinct groups, each modifying different aspects of the original concept. These groups serve as dimensions, allowing the combination of variants from different groups to create new definitions. The study also established a partial ordering of their strength and identified which definitions exhibit properties like composition, post-processing, and convexity. Guingona et al. (2023) compared approximate and probabilistic differential privacy, exploring the implications of their relationship and finding that significant adjustments to one parameter necessitate substantial adjustments to another. Ünsal and Önen (2023) investigated the connection between differential privacy and information-theoretic measures, demonstrating how these connections can provide novel interpretations of privacy in systems employing differential privacy.
Our work contributes to this growing body of knowledge by introducing Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP), and establishing their relationships with Local Differential Privacy (LDP). Specifically, we prove that $\xi$-LDP implies $\xi$-MBP, which in turn implies $2\xi$-LDP. We also demonstrate that MBP implies ABP.
Our approach to reductions differs from previous works in that we focus on Bayesian privacy notions and their connections to LDP. While previous taxonomies have primarily focused on variations of differential privacy, our work bridges the gap between traditional DP concepts and Bayesian privacy notions. This provides a new perspective on privacy guarantees that takes into account an adversary’s prior knowledge and inference capabilities.
The content of our taxonomy complements existing work by:
@@
    Introducing new Bayesian privacy metrics (MBP and ABP)
@@
    Establishing concrete relationships between these Bayesian metrics and LDP
@@
    Providing a bridge between traditional DP concepts and Bayesian privacy notions
These results offer new insights into the design of privacy-preserving algorithms, particularly in scenarios where adversaries may have prior knowledge about the data distribution. By situating our work within the broader context of privacy taxonomy research, we contribute to a more comprehensive understanding of privacy definitions and their interrelationships, aiding researchers and practitioners in selecting appropriate privacy metrics for their specific use cases.
II-C Intersections of Privacy and PAC Learning
The dynamic interplay between privacy and learning theory, with a spotlight on PAC learning, has attracted considerable scholarly inquiry. In groundbreaking research, Kasiviswanathan et al. (2011) scrutinized the concept classes amenable to learning under differential privacy. Their findings revealed that a vast array of classes that are learnable by non-private algorithms can also be mastered privately, albeit with a polynomial increase in sample size. This study highlighted the striking equivalence between local private learning algorithms and statistical query learning, underscoring that privacy constraints do not substantially impede the learning process.
Concurrently, Bun (2020) delved into the computational nuances of private PAC learning, drawing a clear line between online learning and private PAC learning paradigms. Their work shed light on the evolving landscape of integrating privacy considerations into learning theory, unveiling both the formidable challenges and promising opportunities in the pursuit of efficient private learning.
In a parallel vein, Feldman and Xiao (2014) explored the nexus between differential privacy and communication complexity within the realm of learning algorithms. They established a profound equivalence between the sample complexity of differentially private learning and the randomized one-way communication complexity required to evaluate concepts. This pivotal discovery is monumental as it forges a link between two distinct domains, offering valuable insights into the development of private learning algorithms and presenting a novel lens through which to view the sample complexity bounds of learning tasks. In a more recent development, Agarwal and Singh (2017) made significant strides in the context of online learning by crafting differentially private algorithms for online convex optimization. They demonstrated that it is feasible to uphold privacy without markedly sacrificing the regret bounds, thus underscoring the intricate equilibrium that must be struck between privacy assurances and learning efficacy in dynamic environments.
Building upon these foundational intersections between privacy and PAC learning, our work extends the discourse by delving into the robustness implications of privacy guarantees. While previous studies have predominantly focused on sample complexity and learnability, we take a deeper dive into how privacy-preserving mechanisms can enhance robustness in PAC learning scenarios, and vice versa, how robust learning algorithms can be harnessed to design privacy-preserving mechanisms.
By establishing theoretical linkages between our proposed privacy notions—MBP and ABP—and PAC robustness, we offer a fresh perspective on the privacy-robustness interplay. Our findings augment the existing body of literature by shedding light on how privacy guarantees can bolster the resilience of learning algorithms to data perturbations, and how robust learning techniques can be leveraged to safeguard individual privacy.
II-D Robustness Measurements
Measuring the robustness of classifiers, especially neural network-based ones, has been an active research area in recent years. Hendrycks and Dietterich (2019) proposed two new notions for measuring robustness: corruption robustness and perturbation robustness, which consider a classifier’s average-case performance on corruptions $\mathcal{C}$ and perturbations $\xi$, respectively. In contrast, adversarial robustness measures the worst-case performance on small, additive, classifier-tailored perturbations or corruptions.
For neural network classifiers, Bastani et al. (2016) formalized the notion of $(x_{\ast},\epsilon)$-robust and defined point-wise robustness, $\rho(f,x_{\ast})$, as the $L_{\infty}$ distance from $x_{\ast}$ to the nearest adversarial example. They proposed two statistics to estimate these quantities: adversarial frequency and adversarial severity.
In the context of binary classification, Zhang et al. (2019) characterized model robustness using the robust classification error under the threat model of bounded $\epsilon$ perturbation, decomposing the robust error into the sum of natural error and boundary error to characterize the trade-off between accuracy and robustness.
Cohen et al. (2019) introduced the concept of certifiable robustness, where a classifier is considered certifiably robust if, for any input $\mathbf{x}$, one can easily obtain a guarantee that the classifier’s prediction remains constant within some set around $\mathbf{x}$. By employing randomized smoothing, they demonstrated the feasibility of certified defense on the ImageNet dataset.
These works provide valuable insights and tools for quantifying and understanding the robustness of classifiers, paving the way for developing more robust models in the face of various types of perturbations and corruptions.
II-E Relation between Robustness and Privacy
The relationship between privacy and robustness has been widely discussed in the literature (Table III). Dwork and Lei (2009) explored the field of robust statistics and demonstrated that robust statistical estimators provide a strong foundation for developing differentially private estimators. Pinot et al. (2019) proposed a generalized notion of adversarial robustness using the divergence $D_{\lambda}$ of probabilistic distributions and suggested using randomization methods to improve robustness against adversarial samples, discussing the similarities between differential privacy and adversarial sample robustness.
bansal2020extending introduced the concept of robust privacy, which defines the proximity and indistinguishability between two distributions without relying on a specific notion of accuracy or the underlying data distribution. Asi et al. (2023) proposed a black-box transformation method to convert a robust statistical estimator into a private statistical estimator using the $\rho$-smooth inverse sensitivity mechanism, proving that a private estimator ($\epsilon$-DP) can naturally be robust, albeit with some loss of accuracy in certain high-dimensional statistical tasks and increased computational inefficiency due to the transformation.
Lecuyer et al. (2019) introduced the PixelDP model, which provides meaningful certified robustness boundaries and reasonable assessments of model accuracy degradation across various datasets and deep neural networks, offering higher accuracy on larger models and datasets, as well as stronger robustness against p-norm attacks compared to previous authentication defense methods. Hopkins et al. (2023) presented the first black-box reduction from privacy to robustness, capable of producing private estimators with optimal trade-offs among sample complexity, accuracy, and privacy for a wide range of fundamental high-dimensional parameter estimation problems, proving that, under certain conditions, if a statistical estimation method is robust to small perturbations of the input data, it can also protect the privacy of the input data.
Pillutla et al. (2022) explored the trade-offs among privacy, robustness, and communication in federated learning, proposing a robust aggregation method that replaces the weighted arithmetic mean aggregation with a generalized form of the classical multidimensional median. This method protects the update privacy of devices by iteratively invoking secure multi-party computation primitives and bounds the sensitivity to each device’s contribution, with numerical simulations demonstrating its effectiveness in handling data corruption and parameter update corruption.
TABLE III: relation between privacy and robustness.
Reference Privacy Metric Robustness Metric Relation
Dwork and Lei (2009) $(\epsilon,\delta)$-DP M-estimators robust covariance Propose-Test-Release
matrices robust regression paradigm
Lecuyer et al. (2019) $(\epsilon,\delta)$-PixelDP certifiably robust $(\epsilon,\delta)-DP$+certain
condition $\Rightarrow$ robustness
Pinot et al. (2019) Renyi-DP $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ $D_{\mathcal{P}(\mathcal{\gamma})}$-$(\alpha,\epsilon,\gamma)$ robust
robust $\Leftrightarrow$ Renyi-DP
bansal2020extending Flexible Accuracy Laplace Mechanism robustness $\Leftrightarrow$ privacy
Robust Privacy $M_{Lap}^{f,b}$
Pillutla et al. (2022) secure median’s be compatible only with
average oracle breakdown point more communication cost1
Hopkins et al. (2023) $(\epsilon,\delta)$-DP $(\eta,\beta,\alpha)$-robust robustness $\Leftrightarrow$ privacy
Asi et al. (2023) $\epsilon$-DP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness2
@@ II-E Relation between Robustness and Privacy
our work $\xi$-MBP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness
$\epsilon$-ABP $(\tau,\beta,\alpha)$-robust robustness $\Rightarrow$ privacy
@@
    No quantitative relation is provided in this work

### v5

@@ Abstract
The widespread adoption of machine learning necessitates robust privacy protection alongside algorithmic resilience. While Local Differential Privacy (LDP) provides foundational guarantees, sophisticated adversaries with prior knowledge demand more nuanced Bayesian privacy notions, such as Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP), first introduced by [61]. Concurrently, machine learning systems require inherent robustness against data perturbations and adversarial manipulations. This paper systematically investigates the intricate theoretical relationships among LDP, MBP, and ABP. Crucially, we bridge these privacy concepts with algorithmic robustness, particularly within the Probably Approximately Correct (PAC) learning framework. Our work demonstrates that privacy-preserving mechanisms inherently confer PAC robustness. We present key theoretical results, including the formalization of the established LDP-MBP relationship, novel bounds between MBP and ABP, and a proof demonstrating PAC robustness from MBP. Furthermore, we establish a novel theoretical relationship quantifying how privacy leakage directly influences an algorithm’s input robustness. These results provide a unified theoretical framework for understanding and optimizing the privacy-robustness trade-off, paving the way for the development of more secure, trustworthy, and resilient machine learning systems.
keywords:
federated learning , privacy-utility tradeoffs, privacy-preserved machine learning
1 Introduction
The advent of machine learning (ML) has ushered in transformative changes across diverse domains—from healthcare diagnostics to financial modeling and autonomous systems—driven by its unprecedented capability to extract meaningful insights from complex datasets [4, 31]. However, this pervasive reliance on data, particularly sensitive information, raises critical concerns regarding privacy and security [51]. Simultaneously, as ML models are increasingly deployed in dynamic real-world environments, their reliability under input inaccuracies or noise—quantified as algorithmic robustness—has become paramount [50, 55]. Building trustworthy, deployable AI systems necessitates addressing both challenges holistically.
The traditional paradigm of Local Differential Privacy (LDP) provides a cornerstone for privacy preservation, bounding the influence of any individual data point on computational outputs to obscure its presence [21]. Particularly valuable in distributed settings where users perturb data locally, LDP offers strong theoretical guarantees. Yet sophisticated adversaries may exploit additional attack vectors. Recent research introduces Bayesian privacy frameworks—Average Bayesian Privacy (ABP) and Maximum Bayesian Privacy (MBP)—that model adversaries with prior knowledge and Bayesian inference capabilities [61]. These frameworks deliver more comprehensive protection by accounting for realistic adversarial strategies. These critical domains have largely evolved independently of algorithmic robustness research, creating substantial theoretical and practical gaps: (i) a lack of unified principles connecting nuanced privacy guarantees with robustness, and (ii) mechanisms that optimize one property at the expense of the other, while yielding suboptimal trade-offs. Bridging these gaps is essential for establishing trustworthy next-generation AI foundations.
Our work addresses these challenges through fundamental theoretical innovations that establish quantifiable synergies between privacy preservation and algorithmic robustness. Our contributions are threefold:
  • First, we derive the inaugural quantitative relationship between MBP and ABP (Theorem 4.2). This result precisely bounds expected privacy leakage (ABP) as a function of worst-case protection (MBP), demonstrating that strengthening MBP inherently enhances ABP—providing crucial mechanism-design guidance.
  • Second, we formally establish that privacy-preserving mechanisms inherently confer Probably Approximately Correct (PAC) robustness (Theorem 4.3). This profound equivalence proves an $(\epsilon,\beta,\alpha)$-MBP mechanism guarantees $\alpha$-PAC-Robustness, with quantifiable confidence degradation $\propto\epsilon\beta$, thereby bridging privacy and stability domains.
  • Third, we reveal a novel mathematical duality between privacy leakage and input robustness (Theorem 5.5). Our analysis demonstrates that intentional parameter distortion ($\Delta$) from privacy mechanisms directly amplifies adversarial resilience ($\alpha$), enabling principled joint optimization via strategic noise injection.
Collectively, these contributions establish a theoretical foundation for developing ML systems that simultaneously resist privacy reconstruction attacks and maintain performance under data perturbations. Our framework delivers: (1) quantitative conversions among Bayesian privacy metrics, (2) formal privacy-to-robustness implications, and (3) pathways for unified optimization.
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 reviews related work on privacy measurements, relationships among privacy metrics, intersections of privacy and PAC learning, robustness measurements, and the relationship between privacy and robustness. In Section 3, we introduce our privacy attack and defense framework. In Section 4, we define metrics for privacy and robustness, and establish theoretical connections between privacy notions and PAC robustness. We also introduce practical estimation approaches and analyses for MBP and ABP in D and E. Finally, we conclude the paper and discuss future research directions.
@@
Table 1: Table of Notation
@@
2 Related Work
2.1 Privacy Measurements
The rigorous formalization of differential privacy (DP) by [23] established a gold standard for privacy-preserving data analysis, catalyzing extensive research into privacy quantification. Subsequent works have expanded this foundation along several dimensions: computational feasibility, distributional awareness, and practical applicability. [2]’s pioneering work demonstrated how deep neural networks could be trained under constrained privacy budgets by integrating DP-SGD with optimization techniques like Momentum [49] and AdaGrad [19]. This approach balanced the tension between model utility and privacy preservation, enabling privacy-sensitive deep learning applications.
Bayesian perspectives introduced significant advancements in privacy measurement. [52] developed Bayesian Differential Privacy (BDP), incorporating data distribution considerations into the privacy calculus while establishing formal connections to traditional DP guarantees. Information-theoretic approaches emerged as powerful alternatives, with [47] employing total variation distance for privacy measurement while assessing utility through minimum mean-square error, mutual information, and error probability. Similarly, [9] proposed entropy-based metrics for quantifying online user privacy, acknowledging the contextual nature of privacy risks.
The practical challenges of achieving pure DP motivated alternative quantification frameworks. [25] formalized privacy loss through relative entropy (Kullback-Leibler divergence), measuring information leakage via posterior belief updates. While [28] identified limitations in posterior sampling for privacy protection, they demonstrated the effectiveness of Laplace mechanisms in Bayesian inference contexts. Social network privacy inspired novel metrics like Audience and Reachability [5], while [18] pioneered minimax formulations for information leakage using KL divergence.
Our work contributes to this landscape by adopting Jensen-Shannon (JS) divergence for privacy measurement, following [61, 63]. This choice is theoretically grounded in JS divergence’s symmetry and triangle inequality properties [44, 26], which enable more nuanced quantification of privacy-utility trade-offs. This approach provides distinct advantages for analyzing the privacy-robustness relationships central to our framework.
2.2 Relation between Distinct Privacy Measurements
The evolution of differential privacy has spurred development of nuanced privacy metrics with complex interrelationships, as summarized in Table 3. Several significant conceptual frameworks have emerged:
Table 2: Taxonomy of privacy metrics and their theoretical relationships
@@ 2.2 Relation between Distinct Privacy Measurements
[24] Concentrated DP DP $\Rightarrow$ Concentrated DP
[42] Rényi-DP Rényi-DP $\Rightarrow$ $(\epsilon,\delta)$-DP
[43]2 IND-CDP, SIM-CDP SIM-CDP $\Rightarrow$ SIM∀∃-CDP3 $\Rightarrow$ IND-CDP
[17] $f$-DP Gaussian DP4 $\Rightarrow$ Rényi-DP
[10] Distributional Privacy Distributional Privacy $\Rightarrow$ DP
[53] Bayesian-DP Bayesian DP $\Rightarrow$ DP
[14] Mutual-Information DP MI-DP $\Leftrightarrow$ DP (asymptotically)
Our Work $\xi$-MBP, $\epsilon$-ABP MBP $\Rightarrow$ ABP
@@
    All "DP"s denote differential privacy variants
@@
    Introduces Computational Differential Privacy (CDP)
@@
    Gaussian DP is a specific $f$-DP instance
Refinements of DP: [24] introduced Concentrated DP, focusing on mean privacy leakage. [42]’s Rényi-DP unified various DP relaxations through $\alpha$-divergences. Computational perspectives emerged with [43]’s IND-CDP and SIM-CDP, which formalized privacy against bounded adversaries. [17] advanced $f$-DP and Gaussian DP, providing tighter composition bounds.
Distribution-Aware Privacy: [10] proposed Distributional Privacy, offering stronger protection than traditional DP through synthetic data generation. [53] introduced Bayesian DP that incorporates dataset distributions while maintaining DP compatibility. Local DP solutions addressed scenarios without trusted third parties [13], with convergence-privacy trade-offs analyzed by [20] and practical mechanisms developed by [16].
Taxonomic Frameworks: Several studies systematically compared privacy notions. [14] established asymptotic equivalence between Mutual-Information DP and traditional DP. [41] clarified distinctions between approximate and probabilistic DP. [15] classified privacy definitions into seven dimensions with partial ordering. [30] compared approximate and probabilistic DP, while [54] connected DP to information-theoretic measures.
Our contribution introduces Maximum Bayesian Privacy (MBP) and Average Bayesian Privacy (ABP) as novel Bayesian privacy metrics. Crucially, we integrate the established equivalence between Local DP and Maximum Bayesian Privacy from [34] into our framework, while demonstrating new connections to robustness. Specifically:
  • We formalize the known relationship: $\xi$-LDP $\Rightarrow$ $\xi$-MBP $\Rightarrow$ $2\xi$-LDP [34]
  • We prove a novel quantitative relationship: MBP $\Rightarrow$ ABP
This synthesis bridges Bayesian privacy frameworks with traditional DP concepts, providing new perspectives for adversarial settings where prior knowledge affects privacy risks. Our taxonomy complements existing work by: (1) Introducing Bayesian metrics sensitive to prior distributions, (2) Establishing quantitative conversions between Bayesian privacy notions (MBP and ABP), and (3) Enabling principled algorithm design for contexts with informed adversaries.
2.3 Intersections of Privacy and PAC Learning
The convergence of privacy guarantees and Probably Approximately Correct (PAC) learning frameworks has generated profound theoretical insights. Seminal work by [36] established that most concept classes learnable non-privately remain learnable under DP constraints, albeit with polynomial sample complexity increases. Their discovery of equivalence between local private learning and statistical query models demonstrated privacy need not fundamentally limit learnability.
Computational aspects were explored by [11], who delineated connections between private PAC learning and online learning paradigms. [27] revealed a profound equivalence between differentially private sample complexity and randomized one-way communication complexity, bridging two distinct theoretical domains. Practical implementations advanced with [3]’s differentially private online convex optimization that maintained near-optimal regret bounds.
Our work extends this research trajectory by investigating how privacy-preserving mechanisms inherently confer PAC robustness. While prior studies focused primarily on learnability and sample complexity, we establish formal connections between Bayesian privacy frameworks (MBP/ABP) and resilience to data perturbations. This provides new theoretical foundations for developing algorithms that simultaneously ensure: (1) Provable privacy against reconstruction attacks, and (2) Certified robustness against adversarial inputs–dual objectives previously addressed separately.
2.4 Robustness Measurements
Robustness quantification has evolved along two primary dimensions: adversarial resistance and distributional stability. [32] introduced the critical distinction between corruption robustness (average-case performance under natural corruptions) and perturbation robustness (targeted adversarial resistance), framing robustness as a multi-faceted requirement. For neural networks, [8] formalized point-wise robustness through adversarial frequency and severity metrics, enabling granular vulnerability assessment.
Theoretical foundations advanced with [58]’s decomposition of robust classification error into natural and boundary components, elucidating accuracy-robustness trade-offs. Practical certification methods emerged through [12]’s randomized smoothing approach, which scaled certifiable robustness to ImageNet-scale problems. These developments collectively established robustness measurement as a rigorous subfield with both theoretical depth and practical applicability.
Our robustness framework builds upon these foundations while introducing PAC-formalized robustness guarantees that explicitly connect to privacy-preserving mechanisms. This dual perspective enables unified analysis of algorithms providing both privacy and robustness certificates.
2.5 Relation between Robustness and Privacy
The intrinsic connection between algorithmic privacy and robustness has emerged as a fundamental research frontier, with key relationships summarized in Table 3. Early connections were identified by [22], who demonstrated how robust statistical estimators naturally enable differentially private estimation through the Propose-Test-Release paradigm. [46] later established formal equivalences between Rényi-DP and distributional robustness through $D_{\lambda}$ divergences.
Three significant research threads have developed:
@@ 2.5 Relation between Robustness and Privacy
    Privacy-to-Robustness: [39]’s PixelDP framework demonstrated how $(\epsilon,\delta)$-DP mechanisms can provide certified robustness guarantees. [7] introduced robust privacy as a generalized indistinguishability criterion.
@@
    Robustness-to-Privacy: [6] showed robust estimators can be transformed into private estimators via $\rho$-smooth inverse sensitivity mechanisms. [33] established black-box reductions from robustness to privacy with optimal trade-offs.
@@
    Federated Settings: [45] developed median-based aggregation providing simultaneous privacy, robustness, and communication efficiency in federated learning, though with communication trade-offs.
Our work advances all three directions through several key contributions:
@@
  • Introduce PAC-formalized robustness with explicit parameter relationships to privacy guarantees
  • Develop quantitative conversion bounds between privacy and robustness parameters
  • Provide the first unified framework connecting Bayesian privacy with PAC robustness
These contributions offer new theoretical tools for co-designing algorithms that simultaneously optimize privacy and robustness–a critical requirement for trustworthy AI systems in sensitive domains.
Table 3: Theoretical relationships between privacy and robustness frameworks
Reference Privacy Metric Robustness Metric Theoretical Relation
[22] $(\epsilon,\delta)$-DP M-estimators Propose-Test-Release
Robust covariance paradigm
Robust regression
[39] $(\epsilon,\delta)$-PixelDP Certifiable robustness $(\epsilon,\delta)$-DP + conditions
$\Rightarrow$ robustness
[46] Rényi-DP $D_{\mathcal{P}(\gamma)}$-$(\alpha,\epsilon,\gamma)$ robust Equivalence:
$D_{\mathcal{P}(\gamma)}$-robust $\Leftrightarrow$ Rényi-DP
[7] Flexible Accuracy Laplace Mechanism Duality:
Robust Privacy $M_{Lap}^{f,b}$ robustness $\Leftrightarrow$ privacy
[45] Secure aggregation Median breakdown point Compatibility requires
increased communication1
[33] $(\epsilon,\delta)$-DP $(\eta,\beta,\alpha)$-robust Robustness $\Leftrightarrow$ privacy
[6] $\epsilon$-DP $(\tau,\beta,\alpha)$-robust Privacy $\Rightarrow$ robustness2
[…]
Our work $\xi$-MBP $(\tau,\beta,\alpha)$-robust Bidirectional:
$\epsilon$-ABP $(\tau,\beta,\alpha)$-robust privacy $\Leftrightarrow$ robustness (MBP)
robustness $\Rightarrow$ privacy (ABP)
@@
    No quantitative relationship established